Determine how much pure water would have to be mixed with 60ml of 80% sulfuric acid to create a 50% solution of sulfuric acid solution?
Answer by Alan3354(69443)

Determine how much pure water would have to be mixed with 60ml of 80% sulfuric acid to create a 50% solution of sulfuric acid solution?
----------------
60*0.8 = 48 ml of acid.
To be 50%, 48 ml of water are needed --> 96 ml of solution.
96 - 60 = 36 ml of water to add.
===============
Or,
x*0 + 60*80 = (60 + x)*50
3000 + 50x = 4800
50x = 1800
x = 36 ml